in a community of 500 people,the 75th percentile age is 65 years while the 25th percentile age is 15 years. How many of the people are between 15 and 65 years.
Answer by Edwin McCravy(20062)

in a community of 500 people,the 75th percentile age is 65 years
 
That means that 75% of the 500 are younger than 65.

Since 75% of 500 = (0.75)(500) = 375, 375 people are younger than 65

while the 25th percentile age is 15 years.
 
That means that 25% of the 500 are younger than 15.

Since 25% of 500 = (0.25)(500) = 375, 125 people are younger than 15.

How many of the people are between 15 and 65 years.
We just take away the 125 that are younger than 15 from the number that
are younger than 65:

375 - 125 = 250 people between 15 and 65 years.

Edwin